# PROV-AGENT: Unified Provenance for Tracking

### v1

@@ Abstract
Foundation models, such as Large Language Models (LLMs), are increasingly used as core components of AI agents in complex, large-scale workflows across federated and heterogeneous environments. In agentic workflows, autonomous agents plan tasks, interact with humans and peers, and shape scientific outcomes. This makes transparency, traceability, reproducibility, and reliability essential. However, AI-based agents can hallucinate or reason incorrectly, and their decisions may propagate errors through the workflow, especially when one agent’s output feeds into another’s input. Therefore, fine-grained provenance is essential to link agent decisions, their end-to-end context, and downstream impacts. While provenance techniques have long supported reproducibility and workflow data understanding, they fail to capture and relate agent-centric metadata (prompts, responses, and decisions) with the rest of the workflow. In this paper, we introduce PROV-AGENT, a provenance model that extends W3C PROV and leverages the Model Context Protocol (MCP) to integrate agent interactions into end-to-end workflow provenance. Our contributions include: (1) a provenance model tailored for agentic workflows, (2) a near real-time, open-source system for capturing agentic provenance, and (3) a cross-facility evaluation spanning edge, cloud, and HPC environments, demonstrating support for critical provenance queries and agent reliability analysis.
@@ I Introduction
In this paper, we build on foundational efforts in workflow provenance research to introduce a framework that captures both traditional workflow metadata and AI agent interactions. Our contributions are threefold: (1) PROV-AGENT, a provenance model that extends the W3C PROV [7] standard and incorporates concepts from the Model Context Protocol (MCP) [8] to represent agent actions and their connections to data and workflow tasks; (2) an open-source system [9] for capturing agentic provenance during execution; and (3) a preliminary evaluation with a cross-facility agentic workflow involving edge devices, cloud services, and HPC systems.
@@ IV-A Implementation
Rather than building a new provenance system from scratch, we extend Flowcept [9], an open-source distributed provenance framework designed for complex, heterogeneous workflows spanning experimental facilities at the edge, cloud platforms, and HPC environments. Flowcept uses a federated, broker-based model where raw provenance data, which may come in varied formats and schemas, can be streamed from instrumented scripts, observability hooks in workflow tools (e.g., Dask, MLflow), and from data streaming services and storage layers such as Redis, Kafka, SQLite, file systems, and object stores while the workflows run [18]. A central consolidation service unifies, curates this data into a persistent provenance database while applying a W3C PROV-extended model, making it well-suited for capturing and contextualizing provenance in end-to-end agentic workflows.
@@
Tool executions are often informed by one or more AIModel calls. Given our driving use cases and that most agentic workflow users employ LLMs for their agents, this first implementation of PROV-AGENT focuses on supporting LLMs by providing a generic wrapper for abstract LLM objects, compatible with models from popular LLM interfaces, including CrewAI, LangChain, and OpenAI. Whenever a prompt is sent to an LLM service provider in the cloud (e.g., OpenAI, SambaNova, Azure), the wrapper captures the prompt, response, model metadata (e.g., provider, name, and parameters like temperature), and optional telemetry such as response time. Listing 4 shows an MCP tool example annotated with the decorator and using the wrapper FowceptLLM. Model metadata are recorded within an instance of AIModel and each invocation is recorded as an AIModelInvocation activity and linked to the model, prompt, and response, according to the defined relationships. When a tool depends on LLM results, Flowcept establishes a wasInformedBy relationship from the AgentTool to the relevant AIModelInvocation activities. While this implementation records only the agent’s ID and name, the model supports extended metadata, such as model and tools’ version control state, and further configuration parameters.

### v2

@@ Abstract
Large Language Models (LLMs) and other foundation models are increasingly used as the core of AI agents. In agentic workflows, these agents plan tasks, interact with humans and peers, and influence scientific outcomes across federated and heterogeneous workflows. However, agents can hallucinate or reason incorrectly, propagating errors when one agent’s output becomes another’s input. Thus, assuring that agents’ actions are transparent, traceable, reproducible, and reliable is critical to assess hallucination risks and mitigate their workflow impacts. While provenance techniques have long supported these principles, existing methods fail to capture and relate agent-centric metadata such as prompts, responses, and decisions with the broader workflow context and downstream outcomes. In this paper, we introduce PROV-AGENT, a provenance model that extends W3C PROV and leverages the Model Context Protocol (MCP) and data observability to integrate agent interactions into end-to-end workflow provenance. Our contributions include: (1) a provenance model tailored for agentic workflows, (2) a near real-time, open-source system for capturing agentic provenance, and (3) a cross-facility evaluation spanning edge, cloud, and HPC environments, demonstrating support for critical provenance queries and agent reliability analysis.
@@ I Introduction
In this paper, we build on foundational efforts in workflow provenance research to introduce a framework that captures both traditional workflow metadata and AI agent interactions. Our contributions are threefold: (1) PROV-AGENT, a provenance model that extends the W3C PROV [7] standard and incorporates concepts from the Model Context Protocol (MCP) [8] to represent agent actions and their connections to data and workflow tasks; (2) an open-source system [9] for data observability and runtime agentic provenance capture in workflows; and (3) a preliminary evaluation with a cross-facility agentic workflow involving edge devices, cloud services, and HPC systems.
@@ IV-A Implementation
Rather than building a new provenance system from scratch, we extend Flowcept [9], an open-source distributed provenance framework designed for complex, heterogeneous workflows spanning experimental facilities at the edge, cloud platforms, and HPC environments. Flowcept uses a federated, broker-based model where raw provenance data, which may come in varied formats and schemas, can be streamed from instrumented scripts, data observability hooks in workflow tools (e.g., Dask, MLflow), and from data streaming services and storage layers such as Redis, Kafka, SQLite, file systems, and object stores while the workflows run [18]. A central consolidation service unifies, curates this data into a persistent provenance database while applying a W3C PROV-extended model, making it well-suited for capturing and contextualizing provenance in end-to-end agentic workflows.
@@
Tool executions are often informed by one or more AIModel calls. Given our driving use cases and that most agentic workflow users employ LLMs for their agents, this first implementation of PROV-AGENT focuses on supporting LLMs by providing a generic wrapper for abstract LLM objects, compatible with models from popular LLM interfaces, including CrewAI, LangChain, and OpenAI. Whenever a prompt is sent to an LLM service provider in the cloud (e.g., OpenAI, SambaNova, Azure), the wrapper captures the prompt, response, model metadata (e.g., provider, name, and parameters like temperature), and optional telemetry such as response time. Fig. 4 shows an MCP tool example annotated with the decorator and using the wrapper FlowceptLLM. Model metadata are recorded within an instance of AIModel and each invocation is recorded as an AIModelInvocation activity and linked to the model, prompt, and response, according to the defined relationships. When a tool depends on LLM results, Flowcept establishes a wasInformedBy relationship from the AgentTool to the relevant AIModelInvocation activities. While this implementation records only the agent’s ID and name, the model supports extended metadata, such as model and tools’ version control state, and further configuration parameters.